Year Cash Flow 0 −$7,700           1 5,100           2 5,100           3 4,400      &nb

Year Cash Flow
0 −$7,700          
1 5,100          
2 5,100          
3 4,400          

  

Required :
(a) What is the profitability index for the cashflows if the relevant discount rate is 9 percent?
(Click to select)$1.687$1.606$1.526$1.558$1.655

   

(b) What is the profitability index for the cashflows if the relevant discount rate is 17 percent?
(Click to select)$1.449$1.477$1.365$1.407$1.336

   

(c) What is the profitability index for the cashflows if the relevant discount rate is 27 percent?
(Click to select)1.2471.2111.1511.1751.272
Answer #1

The profitability index is defined as the PV of the cash inflows divided by the PV of the cash outflows.

(a) The equation for the profitability index at a required return of 9 percent is:

PI = ($5,100 / 1.09 + $5,100 / 1.092 + $4,400 / 1.093) / $7,700 = 1.606

(b) The equation for the profitability index at a required return of 17 percent is:

PI = ($5,100 / 1.17 + $5,100 / 1.172 + $4,400 / 1.173) / $7,700 = 1.407

(c) The equation for the profitability index at a required return of 27 percent is:

PI = ($5,100 / 1.27 + $5,100 / 1.272 + $4,400 / 1.273) / $7,700 = 1.211